OpenStudy (anonymous):

can you solve this math problem: Solve by substitution method. x + y = 6 y= x - 2

OpenStudy (mathstudent55):

The substitution method is where you solve one equation for one variable, then you plug that in the other equation. In other words, solve one equation for y, for example. Then whatever y is equal to, put that in for y in the other equation. If you look at both equations, the second one is already solved for y.

OpenStudy (mathstudent55):

That means the first step is done and you know y = x - 2. Now where you see y in the other equation, put in x - 2 bec that's what y is equal to.

OpenStudy (mathstudent55):

First equation: x + y = 6 y = x - 2, so first equation becomes: x + x - 2 = 6 Now solve for x

OpenStudy (mathstudent55):

Once you have a value for x, plug that into either one of the two original equations and solve for y.